# for Large Language Models

### v1

@@ Abstract
Machine Learning, ICML
@@ 1 Introduction
Large language models (OpenAI et al., 2024; Gemini, 2024; DeepSeek-AI et al., 2024) have displayed remarkable proficiency across a wide range of tasks, mainly because they excel in instruction following, reasoning, long context understanding, code generation, and so on (Ouyang et al., 2022; Cobbe et al., 2021; Su et al., 2024; Roziere et al., 2023; Lu et al., 2024; Sun et al., 2024). Inherently, these capabilities are language-agnostic. Consider a simple task like the acquisition of mathematical concept: the numerical outcome remains consistent regardless of whether one learns the arithmetic expression $1+1=2$ in English or Chinese. Similarly, when it comes to coding tasks, the choice between English or Chinese for articulating these instructions does not alter the fundamental logic of the code. However, numerous empirical studies have shown that LLMs’ multilingual performance is quite unbalanced across different languages when handling same tasks (Shi et al., 2023; Zhu et al., 2024; Qi et al., 2023).
@@
However, current benchmarks (Hendrycks et al., 2021; Lai et al., 2023; Singh et al., 2024; Wang et al., 2024a) do not support comprehensive testing of the language-agnostic abilities of LLMs, particularly in low-resource language settings, for several reasons. Tasks like XWinograd (Muennighoff et al., 2023) and XStoryCloze (Lin et al., 2022), based on multiple-choice formats, do not fully evaluate the generative capacities of LLMs. Additionally, the limited language overlap across existing benchmarks poses challenges in assessing LLM performance in specific languages. Recently, P-MMEval (Zhang et al., 2024) is proposed as a multilingual multitask benchmark, with the majority of its tasks still following a multiple-choice format. While it includes assessments like MGSM (Shi et al., 2023) and MIFEVAL that cover partial language-agnostic capabilities, LLMs exhibit remarkable performance, as shown in Table 1. This narrow focus leaves a significant gap between research evaluation and real-world applications.
To tackle this problem, we develop a comprehensive, multi-way, and challenging multilingual evaluation suite, called BenchMAX, to help the community better analyze and improve the language-agnostic capabilities of LLMs. Covering 17 languages [^3] The 17 languages include English, Spanish, French, German, Russian, Bengali, Japanese, Thai, Swahili, Chinese, Telugu, Arabic, Korean, Serbian, Czech, Hungarian, and Vietnamese., BenchMAX not only includes a broader range of language families but also emphasizes the diversity of writing systems across languages (Table 1). As demonstrated in Table 1, BenchMAX increases the percentage of studied languages that utilize the non-Latin script.
Tasks Llama3.1 70B Qwen2.5 72B # LG # LG-Family Diversity $R_{\mathrm{non-Latin}}$
@@
Meanwhile, BenchMAX highlights diverse language-agnostic advanced capabilities (Figure 1). We assess instruction following capability with rule-based (Zhou et al., 2023) and model-based (Li et al., 2024) evaluations, code generation capability in diverse scenarios (function-completion (Liu et al., 2024) / problem solving (Jain et al., 2024)), long context understanding capability (Hsieh et al., 2024), a verity of reasoning in math (Shi et al., 2023) and science (Rein et al., 2023), tool use (Srinivasan et al., 2023) in agent environments, and general (Costa-jussà et al., 2022) / domain translation. Domain translation, a by-product of data construction, poses a new challenge for LLM by necessitating fine-grained control and domain-specific terminology understanding over the translation process.
To ensure high quality, we devise an annotation framework to optimize the dataset quality with human effort and LLM feedback. The process involves translating data from English to selected non-English languages using machine translation systems, post-editing each sample by three native-speaking annotators across all tasks, and picking the final translation version using a superior LLM that involves swapping sample positions for debiasing (Wang et al., 2024b; Li et al., 2024).
@@
  • We develop a comprehensive, multi-way multilingual benchmark across 17 languages for evaluating 6 crucial capabilities of LLMs on 10 diverse tasks.
  • We propose a pipeline for curating high-quality mutlilingual datasets incorporating both human annotation and LLM-as-a-judge.
  • Leading multilingual LLMs are evaluted on BenchMAX, and the related analyses provide a further understanding of the language-agnostic capabilities.
2 Related Work
Prior to the era of large language models, most multilingual benchmarks are designed to evaluate discriminative models and take the form of classification tasks, such as XNLI (Conneau et al., 2018), XCOPA (Ponti et al., 2020), XCSQA (Talmor et al., 2019), and so on (Lin et al., 2022; Muennighoff et al., 2023). However, due to their limited task complexity and the lack of diversity in format, these tasks become less practical. Recently, MGSM has become the most frequently used dataset in technical reports from leading LLM teams (Dubey et al., 2024; Gemini, 2024; OpenAI, 2024), which measures the mathematical reasoning capability across eleven languages. In this paper, we extend it to cover six additional diverse languages. Another widely used benchmark is the multilingual translated version of MMLU (Hendrycks et al., 2021; Lai et al., 2023; Singh et al., 2024), which assesses LLMs on knowledge-intensive tasks. However, due to the lack of a unified dataset version, scores are often difficult to compare across studies. Moreover, recent analyses have revealed that MMLU contains numerous ground truth errors (Gema et al., 2024), obscuring the accurate evaluation of LLM capabilities. To address these limitations, our work builds upon GPQA dataset (Rein et al., 2023) instead of MMLU, which offers higher-quality annotations and poses greater challenges in domain-specific knowledge and reasoning evaluation. In addition to curating multilingual versions of MGSM and GPQA, we incorporate a broader range of capabilities, including long context modeling (Hsieh et al., 2024), tool use (Srinivasan et al., 2023), instruction following (Zhou et al., 2023), and more. This allows our benchmark to evaluate LLMs’ multilingual capabilities more comprehensively compared to previous aggregated benchmarks, such as SeaEval (Wang et al., 2024a) and P-MMEval (Zhang et al., 2024). More importantly, all translations except the long context data in our benchmark are post-edited by native human experts after machine translation. This significantly improves both the quality and reliability of the dataset.
Table 2: BenchMAX covers a wide range of language families and script systems.
@@ 2 Related Work
Table 3: Selection of core capabilities and details of task data. For IFEval, we filter out all language-specific instructions, thus remaining 429 samples. For Nexus, we only adopt the standardized_queries subset which contains 318 samples. For general translation datasets, the number of samples may vary in different translation directions, according to the number of parallel samples in TED and WMT24.
@@
Following Model-based Arena-hard 500 Win Rate Problem Solving LiveCodeBench_v4 713
Reasoning Math MGSM 250 Exact Match Translation General Flores+TED+WMT24 $\geq$1012 spBLEU
@@ 3.2 Capabilities Selection
LLMs have demonstrated proficiency in understanding tasks such as text classification, sentiment analysis, and so on. However, their capabilities transcend text understanding, possessing the following intrinsic capabilities:
  • Instruction Following: Following instructions capability is categorized into two distinct tasks based on evaluation paradigms: rule-based and model-based assessment.
@@
  • Reasoning: The capability to reason through intricate scenarios including both math reasoning and natural scientific (physics, chemistry, and biology) reasoning tasks.
@@
  • Code Generation: We primarily consider Python executable code generation in two settings, function completion and programming problem solving.
@@
  • Long Context Modeling: The ability to extract evidence from lengthy documents. We evaluate this capability through question-answering tasks with long documents  (128k tokens).
@@
  • Tool Use: We assess the capability of utilizing tools effectively to correctly select and invoke a single function from multiple available functions based on given user queries.
@@
  • Translation: Translation involves accurately converting text between languages while preserving semantic meaning. Beyond traditional translation tasks, we introduce the Domain Translation task, a by-product of the BenchMAX construction process. This task challenges models to translate specialized terminology and determine whether specific segments should be translated.
@@
Further details on the datasets, sample sizes, and evaluation metrics are provided in Table 3. More detailed information can be found in Appendix A.
Table 4: One example in rule-based instruction following task, which includes complex constraints. First, we enclose these constraints with special symbols and then use a machine translation system to translate from English to the target language. Finally, we reform the case structure by extracting the constraint from machine translation for human post-editing.
@@
[Original Text]: {prompt: Create an ad copy by expanding ”Get 40 miles per gallon on the highway” in the form of a QA with a weird style. Your response should contain less than 8 sentences. Do not include keywords ’mileage’ or ’fuel’ in your response.
@@
instruction_id_list: [’length_constraints: number_sentences’, ’keywords: forbidden_words’]
@@
kwargs: [{’relation’: ’less than’, ’num_sentences’: 8}, {’forbidden_words’: [’mileage’, ’fuel’]}]}
@@
[Translation Input]: Create an ad copy by expanding ”Get 40 miles per gallon on the highway” in the form of a QA with a weird style. Your response should contain less than 8 sentences. Do not include keywords ’$<$b$>$mileage$<$/b$>$’ or ’$<$b$>$fuel$<$/b$>$’ in your response.
[Google Translation Result]: {CJK}UTF8gbsn以风格怪异的问答形式扩展“在高速公路上每加仑行驶 40 英里”来创建广告文案。您的回复应少于 8 个句子。请勿在回复中包含关键字“<<<b>>>里程<<</b>>>”或“<<<b>>>燃料<<</b>>>”。
[Case Reform] { prompt: {CJK}UTF8gbsn以风格怪异的问答形式扩展“在高速公路上每加仑行驶 40 英里”来创建广告文案。您的回复应少于 8 个句子。请勿在回复中包含关键字“里程”或“燃料”。
@@
instruction_id_list: [’length_constraints:number_sentences’, ’keywords:forbidden_words’] kwargs: [{’relation’: ’less than’, ’num_sentences’: 8}, ’forbidden_words’: [{CJK}UTF8gbsn’里程’, ’燃料’]] }
[Human Post-Editing] { ”prompt”: {CJK}UTF8gbsn以一种奇特风格的问答形式展开“在高速公路上每加仑行驶40英里”这句话,创建为一个广告文案。你的回答应该少于8句话。不要在你的回复中包含关键字“里程”或“燃料”。,
@@
instruction_id_list: [’length_constraints:number_sentences’, ’keywords:forbidden_words’] kwargs: [{’relation’: ’less than’, ’num_sentences’: 8}, ’forbidden_words’: [{CJK}UTF8gbsn’里程’, ’燃料’]] }
Figure 2: The construction process of BenchMAX involves three steps: Step 1) translating data from English to non-English; Step 2) post-editing each sample by three human annotators; Step 3) selecting the final translation version.
3.3 Construction
We select between traditional translators such as Google Translate, and LLM-based ones like GPT-4o, depending on whether the task contains extractable constraints. As illustrated in Figure 3, if the task data contains constraints that are hard to extract, we prompt GPT-4o to translate the data and satisfy the constraints. Otherwise, we use Google Translate along with extraction tools. Extraction tools can include methods for extracting translated keywords by enclosing source keywords with special symbols, and for preserving source constraints by replacing constraints with placeholders before translation and restoring them afterwards.
Figure 3: Flow chart illustrating the constraint extraction and preprocessing pipeline in the first step of our benchmark construction.
Taking an example of rule-based instruction following task as an example, as shown in Table 3.2, it requires extra processing to extract constraints from the translated instruction, as they are needed for verification. Inspired by Yuan et al. (2020), we enclose the keywords in the original instruction with special symbols, making them easy to extract from the translated result. As shown in Table 5, we explore various groups of special symbols and different orders and calculate the recall rates of keywords. Comparing to not using special symbols, apply any symbol group can greatly improve the recalls, while combining different symbol groups in multiple rounds can further improve the recalls. We choose Order 1 as it can achieve better results with fewer groups than Order 2. The detailed recall results are in Appendix B.
Table 5: The recall rates using different groups of special symbols
@@ 3.3 Construction
Step 2: Post-editing each sample by three distinct native-speaking annotators in all tasks.
To ensure high-quality dataset, we implement a multi-round annotation and verification process. 1) Each sample is given to three native-speaking annotators who are proficient in English and their native language. Considering the specialized nature of datasets like Science reasoning, annotators are required to hold at least a Bachelor’s degree. 2) Two automatic verifiers - rule-based verifiers and model-based verifiers - are used to assess the quality of human annotation. Rule-based verifiers are used to ensure the satisfaction of constraints for certain tasks, such as the rule-based instruction following task. For model-based verifiers, we utilize the GEMBA-SQM prompt (Kocmi & Federmann, 2023) and employ Qwen2.5-72B-Instruct, a powerful multilingual model, estimating the quality of translations. Along with providing an overall score, the model offers detailed explanations of translation errors as feedback to annotators. Samples failing the rule-based verifier or scoring below a predefined threshold are identified as failed samples and refined in subsequent iterations. Each manually annotated dataset undergoes at least three iterations.
Step 3: Selecting the final translation version by LLMs.
Initially, we ask a fourth annotator, uninvolved in the annotation process, to choose the final version from the results revised by three individuals. Intriguingly, the selection by the fourth annotator exhibited a strong position bias, often favoring the initial annotation. This preference could be attributed to the uniformly high quality of the annotations, resulting in minimal discernible differences among them.
Debiasing for human annotators is costly in terms of both time and finance, because three translations encompass all permutations of six. Consequently, we employ GPT-4o-mini to select the final translation, as it is a powerful and balanced LLM across different languages. In particular, following Li et al. (2024), we adapt the LLM-Judge system instruction (see Appendix D) to suit pairwise translation evaluation. We firstly shuffle the positions of the three translations and conduct two battles to select a final version. In each battle between two translations, we perform two judgments by swapping their positions and determine one winner. The winner of the first two translations then battles against the third translation, determining the final winner.
Table 6: Performance comparison across models on BenchMAX tasks, averaged over 17 languages. Bold numbers indicate the best performance in each column. ”Func Compl.” refers to Function Completion, ”Prob. Solving” to Problem Solving, and ”Multi Func.” to Multiple Functions scenarios where models must select and call one function from multiple options. Note that although DeepSeek-V3 supports 128 context, its API only supports 64K context, and we cannot deploy it on our server. Therefore, we have not evaluated its long-context capabilities yet.

### v2

@@ Abstract
BenchMAX: A Comprehensive Multilingual Evaluation Suite
for Large Language Models
Xu Huang1, Wenhao Zhu1, Hanxu Hu2, Conghui He3, Lei Li4, Shujian Huang1thanks: Corresponding authors, Fei Yuan311footnotemark: 1 1National Key Laboratory for Novel Software Technology, Nanjing University 2University of Zurich,4Carnegie Mellon University 3Shanghai Artificial Intelligence Laboratory {xuhuang,zhuwh}@smail.nju.edu.cn, hanxu.hu@uzh.ch, heconghui@pjlab.org.cn leili@cs.cmu.edu, huangsj@nju.edu.cn, yuanfei@pjlab.org.cn
@@ 1 Introduction
Large Language Models (LLMs;  OpenAI et al., 2024; Gemini, 2024; DeepSeek-AI et al., 2024) have displayed remarkable proficiency across a wide range of tasks, mainly because they excel in instruction following, reasoning, long context understanding, code generation, and so on Ouyang et al. (2022); Cobbe et al. (2021); Su et al. (2024); Roziere et al. (2023); Lu et al. (2024); Sun et al. (2024). Inherently, these capabilities are language-agnostic. Consider a simple task like the acquisition of mathematical concept: the numerical outcome remains consistent regardless of whether one learns the arithmetic expression $1+1=2$ in English or Chinese. Similarly, when it comes to coding tasks, the choice between English or Chinese for articulating these instructions does not alter the fundamental logic of the code. However, numerous empirical studies have shown that LLMs’ multilingual performance is quite unbalanced across different languages when handling same tasks Shi et al. (2023); Zhu et al. (2024); Qi et al. (2023).
@@
However, current benchmarks Hendrycks et al. (2021); Lai et al. (2023); Singh et al. (2024); Wang et al. (2024a) do not support comprehensive testing of the language-agnostic abilities of LLMs, particularly in low-resource language settings, for several reasons. Tasks like XWinograd Muennighoff et al. (2023) and XStoryCloze Lin et al. (2022), based on multiple-choice formats, do not fully evaluate the generative capacities of LLMs. Additionally, the limited language overlap across existing benchmarks poses challenges in assessing LLM performance in diverse languages. Recently, P-MMEval Zhang et al. (2024) is proposed as a multilingual multitask benchmark, with the majority of its tasks still following a multiple-choice format. While it includes assessments like MGSM Shi et al. (2023) and MIFEVAL that cover partial language-agnostic capabilities, LLMs exhibit remarkable performance, as shown in Table 1. This narrow focus leaves a significant gap between research evaluation and real-world applications.
To tackle this problem, we develop a comprehensive, multi-way, and challenging multilingual evaluation suite, called BenchMAX, to help the community better analyze and improve the language-agnostic capabilities of LLMs. Covering 17 languages [^3] The 17 languages include English, Spanish, French, German, Russian, Bengali, Japanese, Thai, Swahili, Chinese, Telugu, Arabic, Korean, Serbian, Czech, Hungarian, and Vietnamese., BenchMAX not only includes a broader range of language families but also emphasizes the diversity of writing systems across languages. As demonstrated in Table 1, BenchMAX increases the percentage of studied languages that utilize the non-Latin script.
Tasks Llama3.1 70B Qwen2.5 72B # LG # LG-Family Diversity $\boldsymbol{R_{\mathrm{non-Latin}}}$
@@
Meanwhile, BenchMAX highlights diverse language-agnostic advanced capabilities (Figure 1). We assess instruction following capability with rule-based Zhou et al. (2023) and model-based Li et al. (2024) evaluations, code generation capability in diverse scenarios (function-completion Liu et al. (2024) / problem solving Jain et al. (2024)), long context understanding capability Hsieh et al. (2024), a verity of reasoning in math Shi et al. (2023) and science Rein et al. (2023), tool use Srinivasan et al. (2023) in agent environments, and general Costa-jussà et al. (2022) / domain translation. Domain translation, a by-product of data construction, poses a new challenge for LLM by necessitating fine-grained control and domain-specific terminology understanding over the translation process.
To ensure high quality, we devise an annotation framework to optimize the dataset quality with human effort and LLM feedback. The process involves translating data from English to selected non-English languages using machine translation systems, post-editing each sample by three native-speaking annotators with multiple iterations across most tasks, and picking the final translation version using a strong LLM that involves swapping sample positions for debiasing Wang et al. (2024b); Li et al. (2024).
@@
  • We develop a comprehensive, multi-way multilingual benchmark across 17 languages for evaluating 6 crucial capabilities on 10 diverse tasks.
  • We propose a pipeline for curating high-quality mutlilingual datasets, involving both human annotation and LLM-as-a-judge.
  • We evaluate leading multilingual LLMs on BenchMAX, and the related analyses provide a further understanding of the language-agnostic capabilities.
@@
Table 2: Besides English, BenchMAX supports 16 non-English languages, covering a wide range of language families and script systems.
@@
Following Model-based m-ArenaHard 500 Win Rate Problem Solving LiveCodeBench_v4 713
Reasoning Math MGSM 250 Exact Match Translation General Flores+TED+WMT24 [[[[1012, 4049]]]] spBLEU
@@
Table 3: Selection of core capabilities and details of task data. For IFEval, we filter out all language-specific instructions, thus remaining 429 samples. For Nexus, we only adopt the standardized_queries subset which contains 318 samples. For general translation datasets, the number of samples may vary in different translation directions, according to the number of parallel samples in TED and WMT24. The datasets of the model-based instruction following task and math reasoning are expanded from existing multilingual datasets, while others are translated from English datasets.
2 Related Work
Prior to the era of Large Language Models  (LLMs;  OpenAI et al., 2024; Gemini, 2024; DeepSeek-AI et al., 2024), most multilingual benchmarks are designed to evaluate discriminative models and take the form of classification tasks, such as XNLI Conneau et al. (2018), XCOPA Ponti et al. (2020), XCSQA Talmor et al. (2019), and so on Lin et al. (2022); Muennighoff et al. (2023). However, due to their limited task complexity and the lack of diversity in format, these tasks become less practical. Recently, MGSM has become the most frequently used dataset in papers and reports from leading LLM teams Dubey et al. (2024); Gemini (2024); OpenAI (2024), which measures the mathematical reasoning capability across eleven languages. In this paper, we extend it to cover six additional diverse languages. Another widely used benchmark is the multilingual translated version of MMLU Hendrycks et al. (2021); Lai et al. (2023); Singh et al. (2024), which assesses LLMs on knowledge-intensive tasks. However, due to the lack of a unified dataset version, scores are often difficult to compare across studies. Moreover, recent analyses have revealed that MMLU contains numerous ground truth errors Gema et al. (2024), obscuring the accurate evaluation. More recently, Include Romanou et al. (2024) is proposed to evaluate the multilingual regional knowledge, hence it’s questions are not language-agnostic. To address these limitations, our work builds upon GPQA dataset Rein et al. (2023) instead of MMLU, which offers higher-quality annotations and poses greater challenges in domain-specific knowledge and reasoning evaluation. In addition to curating multilingual versions of MGSM and GPQA, we incorporate a broader range of capabilities, including long context modeling Hsieh et al. (2024), tool use Srinivasan et al. (2023), instruction following Zhou et al. (2023), and more. This allows our benchmark to evaluate LLMs’ multilingual capabilities more comprehensively compared to previous aggregated benchmarks, such as SeaEval Wang et al. (2024a) and P-MMEval Zhang et al. (2024). More importantly, all translations except the long context data in our benchmark are post-edited by native human experts after machine translation. This significantly improves both the quality and reliability of the dataset.
@@ 3 Benchmark Construction
Figure 2: The construction process involves three steps: Step 1) translating data from English to non-English; Step 2) post-editing each sample by three human annotators; Step 3) selecting the final translation version.
@@ 3.1 Language Selection
Figure 3: Flow chart illustrating the constraint extraction and machine translation pipeline in the first step of our benchmark construction.
@@ 3.2 Capabilities Selection
LLMs have demonstrated proficiency in understanding tasks such as text classification and sentiment analysis, but their capabilities transcend understanding. We construct tasks to evaluate following intrinsic capabilities in multilingual settings:
  • Instruction Following: The capability to follow instructions is evaluated by two distinct tasks with different evaluation paradigms: rule-based and model-based assessment. For the rule-based task, we translate IFEval Zhou et al. (2023) from English to other languages, while we expand m-ArenaHard Dang et al. (2024); Li et al. (2024) to languages we selected for the model-based task.
@@
  • Reasoning: The reasoning capability is assessed through intricate scenarios including math and natural science (physics, chemistry, and biology) problems. We expand MGSM Shi et al. (2023) and GPQA Rein et al. (2023) to 17 languages for the math reasoning and science reasoning tasks.
@@
  • Code Generation: We primarily consider Python code generation in two settings, function completion and programming problem solving. We translate Humaneval+ Liu et al. (2024); Chen et al. (2021) and LiveCodeBench_v4 Jain et al. (2024) from English to other languages.
@@
  • Long Context Modeling: We evaluate the ability to extract evidence from lengthy documents through question-answering tasks with long documents (128k tokens). We build this task based on RULER Hsieh et al. (2024), and translate haystacks, needles, and QA pairs.
@@
  • Tool Use: We assess the ability to correctly select and invoke a single function from multiple available options in response to user queries. We translate the queries in Nexus Srinivasan et al. (2023) to other languages, but leave the APIs in English.
@@
  • Translation: Translation involves accurately converting text between languages while preserving semantic meanings. Beyond traditional translation tasks including Flores, TED, and WMT Costa-jussà et al. (2022); Cettolo et al. (2012); Kocmi et al. (2024), we introduce the Domain Translation task, a by-product of the BenchMAX construction process. This task challenges models to translate specialized terminology and determine whether specific segments should be translated.
@@
The information of the seleted datasets, sample sizes, and evaluation metrics is provided in Table 3. More details can be found in Appendix A.
@@
[Original Text] {prompt: Create an ad copy by expanding “Get 40 miles per gallon on the highway” in the form of a QA with a weird style. Your response should contain less than 8 sentences. Do not include keywords ‘mileage’ or ‘fuel’ in your response.
@@
instruction_id_list: [‘length_constraints: number_sentences’, ‘keywords: forbidden_words’]
@@
kwargs: [{’relation’: ‘less than’, ‘num_sentences’: 8}, {‘forbidden_words’: [‘mileage’, ‘fuel’]}]}
@@
[Translation Input] Create an ad copy by expanding "Get 40 miles per gallon on the highway" in the form of a QA with a weird style. Your response should contain less than 8 sentences. Do not include keywords ‘$<$b$>$mileage$<$/b$>$’ or ‘$<$b$>$fuel$<$/b$>$’ in your response.
[Google Translation Result] {CJK}UTF8gbsn以风格怪异的问答形式扩展“在高速公路上每加仑行驶 40 英里”来创建广告文案。您的回复应少于 8 个句子。请勿在回复中包含关键字“<<<b>>>里程<<</b>>>”或“<<<b>>>燃料<<</b>>>”。
[Postprocessing] {prompt: {CJK}UTF8gbsn以风格怪异的问答形式扩展“在高速公路上每加仑行驶 40 英里”来创建广告文案。您的回复应少于 8 个句子。请勿在回复中包含关键字“里程”或“燃料”。
@@
instruction_id_list: [‘length_constraints: number_sentences’, ‘keywords: forbidden_words’]
@@
kwargs: [{‘relation’: ‘less than’, ‘num_sentences’: 8}, ‘forbidden_words’: [{CJK}UTF8gbsn‘里程’, ‘燃料’]] }
[Human Post-Editing] {prompt: {CJK}UTF8gbsn以一种奇特风格的问答形式展开“在高速公路上每加仑行驶40英里”这句话,创建为一个广告文案。你的回答应该少于8句话。不要在你的回复中包含关键字“里程”或“燃料”。
instruction_id_list: [‘length_constraints: number_sentences’, ‘keywords: forbidden_words’]
kwargs: [{‘relation’: ‘less than’, ‘num_sentences’: 8}, ‘forbidden_words’: [{CJK}UTF8gbsn‘里程’, ‘燃料’]] }
Table 4: One example in rule-based instruction following task, which includes complex constraints. First, we enclose these constraints with special symbols and then translate the prompt from English to the target language by Google Translate. Finally, we postprocess the prompt by extracting constraints into kwargs and removing special symbols for human post-editing.
@@
Table 5: The recall rates of constraints using different groups of special symbols. We choose Order 1, which has fewer steps and produces on-par or better performance than other settings.
3.3 Construction
We select between traditional translators such as Google Translate, and LLM-based ones like GPT-4o, depending on whether a task contains extractable constraints. As illustrated in Figure 3, if the data contains constraints that are hard to extract, we prompt GPT-4o to translate the data and satisfy the constraints. Otherwise, we use Google Translate along with extraction tools. Extraction tools include methods for extracting translated keywords by enclosing source keywords with special symbols, and for preserving source constraints by replacing constraints with placeholders before translation and restoring them afterwards.
Taking an example of rule-based instruction following task as an example, as shown in Table 3.2, it requires extra processing to extract constraints from the translated instruction, as they are needed for verification. Inspired by Yuan et al. (2020), we enclose the keywords in the original instruction with special symbols, making them easy to extract from the translated result. If one symbol fails, another symbol is used to improve recall. As shown in Table 5, we explore various groups of special symbols and different orders, and calculate the recall rates of keywords. Comparing to not using special symbols, apply any symbol group can greatly improve the recalls, while combining different symbol groups in multiple rounds can further improve the recalls. We choose Order 1 as it can achieve better results with fewer groups than Order 2. The detailed recall results are in Appendix B.
Note that in cases where existing multilingual datasets are available, such as MGSM and m-ArenaHard, we extend them to include the supported languages by translating the English data, to minimize additional effort.